Reports 1-1 of 1 Clear search Modify search
yutaro.enomoto - 19:32 Friday 23 February 2018 (4269) Print this report
Sampling frequency of VISETMY model decreased to 2kHz, foton filters messed up and recovered

= Summary =
Changing the sampling frequency of ETMY model made the digital filters of ETMY crazy, and the filters were recovered.

= Detail =
Related to 4261

After some discussion with Miyakawa-san and Yamamoto-san, I editted and ran the ETMY RT model to decrease the sampling frequency, to see if there are any changes in CPU load and ETMY glitch behaviors.
Sampling frequency was changed from 16 kHz to 2 kHz.

After that modification, Okutomi-san pointed out that the degital filters of ETMY got wrong.
The reason is the following:
Foton saves the filters in a time-domain format, so changing the sampling frequency by a factor of 1/8 would result in the change of zero/pole frequencies by a factor of 1/8; foton stores the filter shape in frequency domain only in a comment.

Yamamoto-san kindly tried to fix this messed-up filters but did not succeed, so I had to resotre the filters manually by using the filter information stored in the comment in the foton file.
I have restored almost all the filters. Only one filter, FM10 of MN_DAMP_Y, was not restored and kept blank. --> I will fix it next week

-- Since I recovered them manually, there may be mistakes. If you (and I) find any strange behavior related to the filter of ETMY, please doubt whether the filter is correct.
-- Does anyone know a smart solution for this issue???

Search Help